A bond with a par value of $1,000 and a coupon rate of 7.50% (semiannual coupon) has a yield to maturity of 5.20%. If the bond has 14 years to maturity, what is its current yield?

Group of answer choices

6.11%

6.19%

6.66%

6.24%

0.00%
Answer #1

Current Yield of Bond = Annual Coupon/Current price of bond

M = $1000, n = 14 * 2 = 28 semi-annual periods, C = 7.5% * $1000/2 = $37.50 (semi-annually), i = 5.20%/2 = 2.6% (semi-annually)

P = $739.35 + $487.39

P = $1,226.73

Annual Coupon = 7.5% * $1000 = $75

Current Yield = $75/$1,226.73 = 6.11%